OpenStudy (anonymous):

Change 4x - 3y = -15 to slope-intercept form and state the slope and y-intercept. You must show all work.

OpenStudy (psymon):

So you want to transform what you have to look like y = mx + b. So you have to have y and only y on one side of the equal sign. So, because I like positive #s, Ill add 3y to both sides giving me 4x = -15 + 3y. Next I will add 15 to both sides, giving me 4x + 15 = 3y. Lastly, I get rid of that 3 in front of the y by dividing everything by 3, leaving me with (4x/3) + 5 = y. Now, the slope is the number in front of x and the y intercept is the sole number. So the slope is (4/3) and the y-intercept is 5